# Correlations Are Ruining Your Gradient Descent

### v3

@@ 1 Introduction
In 1998 it was proposed that there might be a perspective beyond typical gradient descent, called natural gradient descent [Amari, 1998], which might overcome some elements of skew and scale in the updates produced by gradient descent. Natural gradient descent has since been explored at the edges of the field of optimization, and deep neural network training [Bernacchia et al., 2018, Martens and Grosse, 2015, Desjardins et al., 2015, Heskes, 2000], with greater stability than traditional second order methods [Dauphin et al., 2014]. However, it is (as a principle) less widely understood, less applied, and less intuitive than it could be.
@@ 2.1 Gradient Descent
The problem which we wish to solve in general is to minimise our loss function and find the optimal set of parameters, effectively to find ${\operatorname{argmin}}_{\bm{\theta}}\,\mathcal{L}(\bm{\theta})$. However, finding this minimum directly is challenging for most interesting problems. Gradient descent proposes a first-order optimisation process by which we identify an update direction (not directly the optimized value) for our parameters based upon a linearization of our loss function. This is often formulated as taking a ‘small step’ in the direction steepest descent of a function, in its gradient direction, such that
@@ 2.2 Natural Gradient Descent
and thus, by finding the minimum of this optimisation (by the first derivative test), we can arrive at a neat formulation of natural gradient descent,
@@
We choose instead to describe this optimiztion in the deterministic regime and provide intuition of its impact, in a manner similar to that of Heskes [2000]. Suppose that, instead of penalizing the Euclidean distance of our parameter change, that we instead penalize the distance travelled in loss space. Mathematically, we are supposing that
@@
It can be shown that by expanding this term with its Taylor series, see Appendix A, one arrives at $\bm{G}=\left\langle\nabla_{\bm{\theta}}\ell^{\top}\nabla_{\bm{\theta}}\ell\right\rangle_{x}$, bringing us to the same solution as found in general for the natural gradients learning rule update, where
${\delta\bm{\theta}}^{\text{NG}}=-{\eta}\left\langle\nabla_{\bm{\theta}}\ell^{\top}\nabla_{\bm{\theta}}\ell\right\rangle^{-1}_{x}\nabla_{\theta}\mathcal{L}.$
@@ 2.3 Natural Gradient Descent for a DNN
Martens and Grosse [2015] (as well as Desjardins et al. [2015] and Bernacchia et al. [2018]) provided a derivation for this quantity. We demonstrate this derivation in Appendix B, and present its conclusions and assumptions here in short. Note that a first assumption is made here, that the natural gradient update can be computed for each layer independently, rather than for the whole network. This approximation not only works in practice [Desjardins et al., 2015] but is also fully theoretically justified for linear networks [Bernacchia et al., 2018]. Taking only a single layer of a network, one may determine that
$\bm{G}_{\bm{W}_{i}}=\langle\nabla_{\theta_{\bm{W}_{i}}}\ell^{\top}\,\nabla_{\bm{\theta}_{\bm{W}_{i}}}\ell\rangle_{\bm{x}_{0}}=\left\langle\text{Vec}\left(\frac{\partial\ell}{\partial\bm{h}_{i}}\bm{x}_{i-1}^{\top}\right)^{\top}\text{Vec}\left(\frac{\partial\ell}{\partial\bm{h}_{i}}\bm{x}_{i-1}^{\top}\right)\right\rangle_{\bm{x}_{0}}\\$
@@
${\delta\bm{\theta}}_{\bm{W}_{i}}=-{\eta}\text{Vec}\left(\left\langle\frac{\partial\ell}{\partial\bm{h}_{i}}^{\top}\frac{\partial\ell}{\partial\bm{h}_{i}}\right\rangle^{-1}_{\bm{y}}\left\langle\frac{\partial\ell}{\partial\bm{h}_{i}}\bm{x}_{i-1}^{\top}\right\rangle_{\bm{x}_{i-1}}\left\langle\bm{x}_{i-1}^{\top}\bm{x}_{i-1}\right\rangle^{-1}_{\bm{x}_{i-1}}\right).\\$
@@ 3 Simple decorrelation mechanisms can rid networks of data correlations
Alternatively, we here describe methods for modifying neural network models such that the neural outputs at each layer are decorrelated via some operation, no matter whether we are doing inference or training. Regular gradient descent in such a case is now closer in its update to natural gradients and furthermore our parameters can now relate in an orthogoal basis.
Note that hereafter we make use of gradient descent to update models and attempt to remove data correlations within our models. Notably, we do not attempt to remove gradient-correlations, the better known aspect of natural gradients. Thus performance is potentially left available via that route.
Existing methods for decorrelation
@@
Few methods have considered iteratively computing a decorrelation matrix directly (i.e. without inversion or matrix decomposition) [Ahmad et al., 2022, Dalm et al., 2024]. The methods which do exist, optimise a matrix for decorrelation alongside the regular weight matrices in a neural network by construction of an appropriate loss function capturing how data should be modified for correlation reduction. These methods, however, were thus far not theoretically linked to natural gradients.
From computational neuroscience however, a number of methods for dynamic and recurrent removal of data correlations were proposed from the perspective of competitive learning and inhibitory control [Földiák, 1990, Pehlevan et al., 2015, Oja, 1989, Vogels et al., 2011]. These methods, in contrast to all the previously described approaches, were however only applied in single layer networks, mostly in unsupervised contexts.
Here we present a method which bridges across all of the above, allowing fast and stable decorrelation while having a formulation within a recurrent system of dynamics.
A note on decorrelation vs whitening
In this work we limit our emphasis on whitening and focus more upon decorrelation. The reason for this is two-fold. First, the off-diagonal elements of the decorrelation matrices cause the greatest impact in skewing gradient descent. The diagonal elements simply scale up/down the gradient vector and this can be trivially dealt with by modern optimizers (e.g. with the Adam optimizer [Kingma and Ba, 2014]). Second, for data in which features are often zero or extremely sparse, normalizing for unit variance can result in extremely large valued activations (distributions with extremely long tails) or unstable updates in the stochastic updating regime.
A novel decorrelation mechanism
Figure 3: Left, existing methods for decorrelation through a system of dynamics (example of Földiák [1990], representative of Pehlevan et al. [2015], Vogels et al. [2011] and more) reduce activity correlations in a manner which is affected by the present scale (eigenvalues) of the decorrelating matrix. Right, existing methods also reduce the scale of all activity in a layer while decorrelating - tending toward the trivial solution of zero activity. The approach proposed in this work solves both of these issues.
@@
    Allows formulation of decorrelation as a distributed dynamical system
We propose that, at all layers of a neural network, one may construct a linear decorrelating transformation of the data, such that $\bar{\bm{x}}_{i}=\bm{M}_{i}\bm{x}_{i}$ where $\bar{\bm{x}}_{i}$ is intended to be a decorrelated form of the data ${\bm{x}}_{i}$. This means that a network now has one additional linear transformation per layer and the full network’s computation is now written
@@
$\bm{M}_{i}\leftarrow g_{i}(\bm{M}_{i}-\eta_{M}\langle\bar{\bm{x}}_{i}\bar{\bm{x}}_{i}^{\top}\rangle\bm{M}_{i})$
[…]
@@
One might enquire as to why this method is useful or interesting. Examining Figure 3, this decorrelation rule is effective at reducing the level of correlation at a network layer regardless of the existing scale (eigenvalues) of the decorrelating matrix - a problem faced by the existing learning rules of Földiák [1990], but also by the similar rules proposed by Pehlevan et al. [2015], Oja [1989], Vogels et al. [2011]. Furthermore, this rule, via the (layer-wise) gain factors $g_{i}$, ensures that the scale (norm) of activities at any given layer remain at the existing scale, rather than reducing, see Figure 3 right. This ensures that decorrelating data does not tend towards the trivial decorrelation solution in which unit activities tend to zero.
@@
Finally, aside from each of these benefits, this system can be easily converted to a system of recurrent dynamics with local-information, see Appendix D. This is particularly a benefit if one wished to implement this method in a neuromorphic device or to link it to biological nervous systems.
@@ 4 Approximate methods for gradient descent work significantly better when coupled with decorrelation
Figure 4: Adding a decorrelation mechanism at every layer of a neural network can massively speed up training convergence speed. Results are shown for backpropagation, feedback alignment, and node perturbation train upon the CIFAR10 in four hidden-layer dense networks, and for backpropagation and feedback alignment upon the CIFAR100 and TinyImageNet classification tasks in five hidden-layer convolutional neural networks. All learning methods were combined with the Adam optimizer [Kingma and Ba, 2014] and the categorical cross entropy loss. Network architectures and training hyperparameters are described in detail in Appendix F. Envelopes show the max and min accuracy levels across five randomly seeded networks. Parameters were selected based upon an independent grid search over decorrelation learning rates and forward learning rates for maximum validation set performance.
@@
Here we show that not only does decorrelation improve training of multi-layer neural networks with the node-perturbation algorithm (as has already been shown by Dalm et al. [2023]) but also significantly improves multi-layer network training when combined with implementation of the feedback alignment algorithm. Networks parameters in pseudocode, as well as the hyperparameters which were used for training are presented in Appendix F and an example of the computational pseudocode provided in Appendix E. Code to reproduce these simulations is available at https://github.com/nasiryahm/CorrelationsRuinGD. Note that all simulations shown below are presented based upon a parameter grid search which was carried out individually for each credit assignment method and best parameters selected for each curve based upon a validation set.

### v4

@@ 1 Introduction
In 1998 it was proposed that there might be a perspective beyond typical gradient descent, called natural gradient descent [Amari, 1998], which might overcome some elements of skew and scale in the updates produced by gradient descent. Natural gradient descent has since been explored at the edges of the field of optimization, and deep neural network training [Bernacchia et al., 2018, Martens & Grosse, 2015, Desjardins et al., 2015, Heskes, 2000], with sometimes greater stability than traditional second order methods [Dauphin et al., 2014], though recently developed second order methods show significant promise [Gupta et al., 2018, Ren & Goldfarb, 2021, Vyas et al., 2024]. Regardless, the principles of natural gradients are less widely understood, less applied, and less intuitive than they could be.
@@ 2.1 Gradient Descent
The problem which we wish to solve in general is to minimise our loss function and find the optimal set of parameters, effectively to find ${\operatorname{argmin}}_{\bm{\theta}}\,\mathcal{L}(\bm{\theta})$. However, finding this minimum directly is challenging for most interesting problems. Gradient descent proposes a first-order optimization process by which we identify an update direction (not directly the optimized value) for our parameters based upon a linearization of our loss function. This is often formulated as taking a ‘small step’ in the direction steepest descent of a function, in its gradient direction, such that
@@ 2.2 Natural Gradient Descent
and thus, by finding the minimum of this optimization (by the first derivative test), we can arrive at a neat formulation of natural gradient descent,
@@
We choose instead to describe this optimization in the deterministic regime and provide intuition of its impact, in a manner similar to that of Heskes [2000]. Suppose that, instead of penalizing the Euclidean distance of our parameter change, that we instead penalize the distance traveled in loss space. Mathematically, we are supposing that
@@
It can be shown that by expanding this term with its Taylor series, see Appendix A, one arrives at $\bm{G}=\left\langle\nabla_{\bm{\theta}}\ell\;\nabla_{\bm{\theta}}\ell^{\top}\right\rangle_{x,y}$, bringing us to the same solution as found in general for the natural gradients learning rule update, where
${\delta\bm{\theta}}^{\text{NG}}=-{\eta}\left\langle\nabla_{\bm{\theta}}\ell\;\nabla_{\bm{\theta}}\ell^{\top}\right\rangle^{-1}_{x,y}\nabla_{\theta}\mathcal{L}.$
@@ 2.3 Natural Gradient Descent for a DNN
Martens & Grosse [2015] (as well as Desjardins et al. [2015] and Bernacchia et al. [2018]) provided a derivation for this quantity. We demonstrate this derivation in Appendix B, and present its conclusions and assumptions here in short. Note that a first assumption is made here, that the natural gradient update can be computed for each layer independently, rather than for the whole network. This approximation not only works in practice [Desjardins et al., 2015] but is also fully theoretically justified for linear networks [Bernacchia et al., 2018]. Taking only a single layer of a network, one may determine that
$\bm{G}_{\bm{W}_{i}}=\langle\nabla_{\theta_{\bm{W}_{i}}}\ell\;\nabla_{\bm{\theta}_{\bm{W}_{i}}}\ell^{\top}\rangle_{\bm{x}_{0},\bm{y}}=\left\langle\text{Vec}\left(\frac{\partial\ell}{\partial\bm{h}_{i}}\bm{x}_{i-1}^{\top}\right)\;\text{Vec}\left(\frac{\partial\ell}{\partial\bm{h}_{i}}\bm{x}_{i-1}^{\top}\right)^{\top}\right\rangle_{\bm{x}_{0},\bm{y}}\\$
@@
${\delta\bm{\theta}}^{\text{NG}}_{\bm{W}_{i}}=-{\eta}\text{Vec}\left(\left\langle\frac{\partial\ell}{\partial\bm{h}_{i}}\frac{\partial\ell}{\partial\bm{h}_{i}}^{\top}\right\rangle^{-1}_{\bm{y}}\left\langle\frac{\partial\ell}{\partial\bm{h}_{i}}\bm{x}_{i-1}^{\top}\right\rangle_{\bm{y},\bm{x}_{i-1}}\left\langle\bm{x}_{i-1}\bm{x}_{i-1}^{\top}\right\rangle^{-1}_{\bm{x}_{i-1}}\right).\\$ (1)
@@ 3 Simple decorrelation mechanisms can rid networks of data correlations
Alternatively, we here describe methods for modifying neural network models such that the neural outputs at each layer are decorrelated via some operation, no matter whether we are doing inference or training. Regular gradient descent in such a case is now closer in its update to natural gradients and furthermore our parameters can now relate in an orthogonal basis. This is explained in greater detail in Section 3.3.
Note that hereafter we make use of gradient descent to update models and attempt to remove data correlations within our models. Notably, we do not attempt to remove gradient correlations, the better known aspect of natural gradients (left most component of Equation 1). Thus performance is potentially left available via that route.
3.1 Existing methods for decorrelation
@@
Few methods have considered iteratively computing a decorrelation matrix directly (i.e. without inversion or matrix decomposition) [Ahmad et al., 2022, Dalm et al., 2024, Sussillo & Abbott, 2009]. Some methods optimise a matrix for decorrelation alongside the regular weight matrices in a neural network by construction of an appropriate loss function capturing how data should be modified for correlation reduction [Ahmad et al., 2022, Dalm et al., 2024]. Other methods [Sussillo & Abbott, 2009] instead do not carry out any decorrelation within a network but instead store a disconnected matrix containing an interatively updated inverse correlation matrix and use this for parameter updating with the goal of achieving recursive least squares optimization. Regardless, these methods were thus far not theoretically linked to natural gradients.
From computational neuroscience however, a number of methods for dynamic and recurrent removal of data correlations have been proposed from the perspective of competitive learning and inhibitory control [Földiák, 1990, Pehlevan et al., 2015, Oja, 1989, Vogels et al., 2011]. Földiák [1990], Pehlevan et al. [2015], Oja [1989], Vogels et al. [2011] describe, in work spanning almost two decades, a set of learning rules between nodes which, via linear recurrent dynamics, push neural activities toward decorrelated states at fixed points of these systems. The rules proposed by all four examples rely upon a simple updating scheme in which recurrent connections within populations of nodes are updated by an ‘anti-hebbian’ parameter update, in short with parameter gradients proportional to node-output correlations. They each, however, contribute a unique perspective on how such an update can be useful, from iterative learning of PCA dimensions [Földiák, 1990, Oja, 1989], through alternative subspace constructions such as multi-dimensional scaling (MDS) extraction [Pehlevan et al., 2015], all the way to an implementation in spiking neural networks which explains the development of real inhibitory synaptic connection structures in neurons [Vogels et al., 2011]. Investigation into even more detailed methods to describe decorrelational inhibitory dynamics continues into contemporary work Lipshutz & Simoncelli [2024]. These methods all provide neat and easily implementable dynamical systems for competition and decorrelation, however in contrast to the machine learning examples above, these methods have all been developed in the context of single layer networks, for unsupervised learning purposes, with little consideration for efficient implementation or application to gradient descent.
Here we present a method which bridges across all of the above, allowing fast and stable decorrelation while having two equivalent formulations: one via a single weight matrix multiplication, and another as the fixed-point of a recurrent system of dynamics.
3.2 A novel decorrelation mechanism
Figure 3: Left, our proposed method can be equivalently represented as either a single linear transformation or the fixed point of a recurrent set of dynamics (see the main text). Middle, existing methods for decorrelation through a system of dynamics (example of Földiák [1990] but representative of Vogels et al. [2011] and more) reduce activity correlations in a manner which is affected by the present scale (eigenvalues) of the decorrelating matrix. Right, existing methods also reduce the scale of all activity in a layer while decorrelating - tending toward the trivial solution of zero activity variance. The approach proposed in this work solves both of these issues.
@@ 3.2 A novel decorrelation mechanism
    Allows formulation of decorrelation as either an efficient linear transformation or equivalently via a distributed dynamical system
We propose that, at all layers of a neural network, one may construct a decorrelating transformation of the data. Efficiently, one can construct this as a linear transformation such that $\bar{\bm{x}}_{i}=\bm{M}_{i}\bm{x}_{i}$ where $\bar{\bm{x}}_{i}$ is intended to be a decorrelated form of the data ${\bm{x}}_{i}$. This means that a network now has one additional linear transformation per layer and the full network’s computation is now written
@@
$\bm{M}_{i}\leftarrow\bm{g}_{i}\circ(\bm{M}_{i}-\eta_{M}\langle\bar{\bm{x}}_{i}\bar{\bm{x}}_{i}^{\top}\rangle\bm{M}_{i})$
where $\eta_{M}$ is a learning rate for this update, and $\bm{g}_{i}$ is a scalar gain to ensures that each of the layer’s decorrelated outputs has the same norm as it had prior to decorrelation at a layer or node level, $\bm{g}_{i}=\langle\bm{x}_{i}^{2}\rangle/\langle\bar{\bm{x}}_{i}^{2}\rangle$. Note that $\circ$ here indicates a Hadamard product (multiplication of each row of the matrix which follows). Note that we find results to be qualitatively independent of whether this normalization occurs at the node or layer-level, but the node-level description has less requirement for layer-wide information sharing and therefore greater biological plausibility.
This update minimises a loss capturing the correlations in $\bar{\bm{x}}$ in the style of Ahmad et al. [2022]. Updates can be carried out in a stochastic fashion by simply measuring the correlations in $\bar{\bm{x}}_{i}$ within a minibatch for iterative updating, see Appendix E for the full pseudocode for such updating. Note that one can further improve upon the level of decorrelation by de-meaning the input data prior to this decorrelation process, such that $\hat{\bm{x}_{i}}=\bm{M}_{i}(\bm{x}_{i}-\bm{\mu}_{i})$ where $\bm{\mu}_{i}$ is a unit-wise learned mean or batch-wise computed mean. We find this to further improve performance in practice.
@@
One might enquire as to why this method is useful or interesting. Examining Figure 3, this decorrelation rule is effective at reducing the level of correlation at a network layer regardless of the existing scale (eigenvalues) of the decorrelating matrix - a problem faced by the existing learning rules of Földiák [1990], but also by the similar rules proposed by Pehlevan et al. [2015], Oja [1989], Vogels et al. [2011]. Furthermore, this rule, via the gain factors $\bm{g}_{i}$, ensures that the scale (norm) of activities at any given layer remain at the existing scale, rather than reducing, see Figure 3 right. This ensures that decorrelating data does not tend towards the trivial decorrelation solution in which unit activities tend to zero.
@@
Finally, aside from each of these benefits, this system can be easily converted to a system of recurrent dynamics with local-information, such that we may also arrive at a decorrelated state by defining,
$\frac{d\bar{\bm{x}}_{i}}{dt}=\bar{\bm{x}}_{i}-\bm{R}_{i}\bm{x}_{i},$
with exact equivalence to our linear decorrelation matrix above, when $\bm{R}_{i}={\bm{M}_{i}}^{-1}$. Notably, the matrix $\bm{R}_{i}$ can be locally updated to match ${\bm{M}_{i}}^{-1}$ via a Shermann-Morrison inverse computation. See Appendix D for a more complete description of such updating. This is particularly a benefit if one wishes to implement our proposed method for modelling of biological nervous systems or for application or translation of our models to analogure/neuromorphic devices.
One consideration to be made when adding any form of decorrelation (or alternative network modification) is it’s additional computational complexity. Our linear transformation-based method for decorrelation adds a square decorrelating matrix for every layer of a neuron network of order. This translate to an additional matrix multiplication of order $\mathcal{O}(n_{i}^{2})$ during inference at each layer, where $n_{i}$ are the number of nodes in layer $i$. Atop this, it adds a corresponding cost to training. Thus, for networks which are particularly wide (rather than narrow and deep), this can have a significant impact upon wall-clock execution time.
Our recurrent formulation requires numerical integration or other form of solving for states to reach their fixed point. Therefore its efficiency is highly dependent upon the exact software/hardware implementation and one should not consider it efficient for applications unless applied in a custom neuromorphic, analogue, or other exotic hardware solution (e.g. nervous systems). The exact computational cost of decorrelation is, for all of these reasons, highly dependent upon network architecture as well as implementation. Therefore it’s utility must be examined on a case-by-case basis.
3.3 Decorrelation better aligns gradient descent with natural gradients
Having motivated and proposed our decorrelation methods, we here briefly demonstrate the impact that decorrelation of data has upon the natural gradients update.
As demonstrated in Section 2.3, the natural gradient update rule for a deep neural network can be expressed in the form of Equation 1. However, in the case in which decorrelation is successful, we have replaced states of layer $i$ such that $\bm{h}_{i}=\bm{W}_{i}\bar{\bm{x}}_{i-1}=\bm{W}_{i}\bm{M}_{i-1}\bm{x}_{i-1}$, and $\langle\bar{\bm{x}}_{i-1}\bar{\bm{x}}_{i-1}^{\top}\rangle=\text{diag}(\langle\bar{\bm{x}}_{i-1}^{2}\rangle)$. As such, the natural gradients update for a decorrelated input state are now computable as
${\delta\bm{\theta}}^{\text{NG-decor}}_{\bm{W}_{i}}=-{\eta}\text{Vec}\left(\left\langle\frac{\partial\ell}{\partial\bm{h}_{i}}\frac{\partial\ell}{\partial\bm{h}_{i}}^{\top}\right\rangle^{-1}_{\bm{y}}\left\langle\frac{\partial\ell}{\partial\bm{h}_{i}}\bar{\bm{x}}_{i-1}^{\top}\right\rangle_{\bm{y},\bar{\bm{x}}_{i-1}}\text{diag}(\langle\bar{\bm{x}}_{i-1}^{2}\rangle)^{-1}\right),\\$
where now our gradient descent update (center) is now modified by a diagonal matrix via right multiplication rather than a full dense matrix. This diagonal matrix can no longer rotate the gradient vector, and thus only has a column-wise re-scaling effect upon the weight matrix update. See Appendix C for a discussion on why we focus on why we focus on decorrelation rather than whitening (which would make this term identity).
In this manner, decorrelation of input states at every layer of a network alleviates one half of the difference between the natural gradient update and regular gradient descent. This is particularly useful for application to algorithms which attempt approximate gradient descent as the gradient signal is compromised but the data signal is not. Therefore we cannot necessarily alleviate gradient correlations, as these are somewhat uncertain, but data correlations can be robustly removed to bring regular gradient descent closer to natural gradients.
@@ 4 Approximate methods for gradient descent work significantly better when coupled with decorrelation
Figure 4: Adding a decorrelation mechanism at every layer of a neural network can massively speed up training convergence speed. Results are shown for backpropagation, feedback alignment, and node perturbation train upon the CIFAR10 in four hidden-layer dense networks, and for backpropagation and feedback alignment upon the CIFAR100 and TinyImageNet classification tasks in five hidden-layer convolutional neural networks. All learning methods were combined with the Adam optimizer [Kingma & Ba, 2014] and the categorical cross entropy loss. Network architectures and training hyperparameters are described in detail in Appendix F. Envelopes show the max and min accuracy levels across five randomly seeded networks.
@@
Here we show that not only does decorrelation improve training of multi-layer neural networks with the node-perturbation algorithm (as has already been shown by Dalm et al. [2023]) but also significantly improves multi-layer network training when combined with implementation of the feedback alignment algorithm. Networks parameters in pseudocode, as well as the hyperparameters which were used for training are presented in Appendix F and an example of the computational pseudocode provided in Appendix E. Note that all simulations shown below are presented based upon a parameter grid search which was carried out individually for each credit assignment method and best parameters selected for each curve based upon a validation set.